What pressure in kPa will 2g of argon exert when placed in a 500mL cylinder at 18oC?
1
Expert's answer
2020-09-29T06:12:30-0400

using ideal gas equation:

PV = nRT

P = "\frac{nRT}{V}"

where n = no. of moles = ("\frac{2}{40}" ) = 0.05moles

Volume = 0.5 L

Temperature = 18oC = 291K


P = "\frac{0.05\times 0.0821 \times 291}{0.5}"

P = 2.38 atm

1 atm = 101.325 kPa

2.38 atm = 241.1535 kPa


Pressure = 241.1535 kPa